Help please and circle the answer and make it clear and visible, thanks. Ive also included a similar problem so you know what the answer to the blank problem should look like

Suppose that the value of the Inventory at Fldos Pet Supply decreases, or depreciates, with time t, in months, where Vis givVſt) = 30- 157 (t+3)2 Suppose that the value of the Inventory at Fldos Pet Supply decreases, or depreciates, with time t, inVit 70- Suppose that the value of the inventory at Fidos Pet Supply CB or deprecates with me, in months where Vis given by t362 VIHE 70 - Suppose that the value of the inventory at Fidos Pet Supply CB or deprecates with me, in months where Vis give
Answer #1

soll V (t) 30 15 t² (t+3) 2 0 9) vlo) 30 24:14 v(s= 30 -15 (542 ( 15/20/? v (30) 18.66 = 30- (23)? V (70) 30 16-21 - 15(70) (2 vet) 70 – 35+ At+3) 2 a ) Vlo) = 70 M(s) 56.33 v (30) = 41.07 v (70) = 37.82 maximumum Valuu 7o y ucts 70 ) graph option B